OpenStudy (anonymous):

(cot^2x/(cscx + 1)) = cos^2x/(sin x(1+sin x)) help me prove the identity simplifying the right side

OpenStudy (anonymous):

you have \[\frac{\cos^2 x}{\sin x (1 + \sin x)}\] first express cos^2 x in terms of sin..you know the pythagorean identity yes?

OpenStudy (anonymous):

not well

OpenStudy (anonymous):

here's a hint: \[\sin^2 x + \cos^2 x = 1\]

OpenStudy (anonymous):

ok so cos^2x = 1 - sin^2x

OpenStudy (anonymous):

right. so if you put that in your equation \[\frac{1 - \sin^2 x}{\sin x (1 + \sin x)}\] do you agree that the numerator is a difference of squares?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if you expand the numerator, what would it be?

OpenStudy (anonymous):

(1 + sin x)( 1 - sin x)

OpenStudy (anonymous):

right. \[\frac{(1 + \sin x)(1 - \sin x)}{\sin x (1 + \sin x)}\] something can be cancelled here

OpenStudy (anonymous):

(1- sin x)/ sin x

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

right. and since the numerator is subtraction, you can separate the terms \[\frac 1{\sin x} - \frac{\sin x}{\sin x}\] simplify

OpenStudy (anonymous):

wait how can you simplify that?

OpenStudy (anonymous):

do you mean how i got there? or how to go from there?

OpenStudy (anonymous):

how to go from there

OpenStudy (anonymous):

it helps if you know the trigonometric identity \(\csc x = \frac 1{\sin x}\) the second term...i assume you know how to simplify that?

OpenStudy (anonymous):

ok so csc x - 1

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

now... a trigonometry magic.... multiply numerator and denominator by csc x + 1

OpenStudy (anonymous):

\[\csc x - 1 \times \frac{\csc x + 1}{\csc x + 1}\]

OpenStudy (anonymous):

so csc^2x + 1/ csc x + 1

OpenStudy (anonymous):

i mean csc^2x - 1 which equals tan^2x

OpenStudy (anonymous):

csc^2 x - 1 is cot^2 x not tan^2 x

OpenStudy (anonymous):

tan^2 x is sec^2 x - 1

OpenStudy (anonymous):

so your equation has been proven. QED

OpenStudy (anonymous):

ok thankyou

OpenStudy (loser66):

LHS, numerator \[cot^2 x= \frac{cos ^2x}{sin^2x}\] denominator \[csc x +1= \frac{1}{sin x}+1= \frac{1+sinx}{sinx}\]combine them \[\huge\frac{\frac{cos^2x}{sin^2x}}{\frac{1+sinx}{sinx}}\] \[= \frac{cos^2x}{sin^2x}*\frac{sinx}{1+sinx}\] cancel out sinx from the terms \[\frac{cos^2x}{sinx(1+sinx)}\] and that's it